Problem 7: Let X and Y be two jointly continuous random variables with joint PDF 4 (x y) otherwise a) Find P(0< Y< 1/2 I x-2) b) For what value of A is it true that P(0 < Y < ½ |X> A)-5/16
Answer #1

a)

f(Y|X = x) = f(x,y) / f(x)

f(X)=\int _0^1\frac{\frac{4}{5}\left(x+y\right)}{x^3}dy=\frac{4x+2}{5x^3}

f(Y|X = x) = 4/5 (x+y) /x^3 / ((4x+2)/5x^3 )

= 4 (x+y)/(4x+2)

f(Y | X = 2) = 4(2 +y)/10) = 0.4 * (y+5)

P(0 <Y< 1/2 |X = 2) =

\int _0^{\frac{1}{2}}0.4\left(y+2\right)dy=0.45

b)

P( 0 < Y <1/2 |X> A) = 5/16

P(0 < Y < 1/2 and X > A) / P(X > A) = 5/16

P(0 < Y < 1/2 and X > A) =

\int _a^{\infty \:}\int _0^{\frac{1}{2}}\frac{\frac{4}{5}\left(x+y\right)}{x^3}dydx=\frac{8a+1}{20a^2}

P(X>A) =

\int _A^{\infty \:}\frac{4x+2}{5x^3}dx=\frac{4A+1}{5A^2}

hence

P( 0 < Y <1/2 |X> A) = 5/16

((8a+ 1)/20a^2 ) /((4A + 1)/5A^2 ) = 5/16

a = 1/12

Please rate